OpenStudy (anonymous):

A lighting rig is a large bank of lights that creates lighting effects at concerts and sporting events. A company makes rigs with lights arranged in polygonal patterns. For example, a triangular rig may have 1, 3, 6, 10, or 15 lights, depending on its size. Which of the following would be an alternate way to represent the total number of lights g(n) in a triangular lighting rig, if n = the number of rows of lights? a. g(n) = g(n-1) + n b. g(g(n)) = g(n) c. g(n) = g(n -1) + 1 d. g(n) = g(n + 1)

OpenStudy (anonymous):

first one i think, we can check it

OpenStudy (anonymous):

\(g(2)=g(1)+2=1+2=3\) yes \(g(3)=g(2)+3=3+3=6\) yes